OpenStudy (technodynamic):

Alright this one is a bit tricky. Please explain this step by step. f(x)=x^2+1 ;Use f(x)+f(2)

OpenStudy (anonymous):

is it the answer is \[[x^{2} + 6]\]

OpenStudy (technodynamic):

can you explain how I should solve this step-by-step?

OpenStudy (anonymous):

\[f(x) = x ^{2} +1\] \[f(x) + f(2) = \left( x ^{2}+1 \right) + \left( 2^{2} +1 \right)\] = x²+1+4+1 = x²+6